### v2

# Covering the permutohedron by affine hyperplanes

Abstract. An almost cover of a finite set in the affine space is a collection of hyperplanes that together cover all points of the set except one. Using the polynomial method we determine the minimum size of an almost cover of the vertex set of the permutohedron and address a few related questions.
1. Introduction
Consider those points in the $n$-dimensional space whose coordinates form a permutation of the first $n$ positive integers. The elements of this set $P_{n}$ are the vertices of a convex $(n-1)$-dimensional polytope called the permutohedron $\Pi_{n-1}$. For $n=3$ it is a regular hexagon, for $n=4$ a truncated octahedron. This polytope has many fascinating properties and can be used to illustrate various concepts in geometry, combinatorics and group theory. For example, it is a simple polytope which tiles the affine hyperplane it is living in. It is also a zonotope: a Minkowski-sum of ($n\choose 2$ line segments. It has $22^{n}-$ 2 facets, corresponding to the nonempty proper subsets of $[n]$. In general, the number of faces of fixed dimension relates to Stirling numbers of the 2nd kind. The symmetric group $S_{n}$ acts on $\Pi_{n-1}$ by permuting the coordinates and can be viewed as a subgroup of the group of its isometries. Using the natural labeling of the vertices by the elements of $S_{n}$ one can identify its edge skeleton with the undirected Cayley graph of $S_{n}$ defined by $n-1$ transpositions. An alternative labeling of the vertices obtained by the anti-isomorphism $\pi\to\pi^{-1}$ reflects the construction of $\Pi_{n-1}$ as the monotone path polytope of the $n$-cube, see [12].
An almost cover of a finite set in the affine space is a collection of hyperplanes that together cover all points of the set except one. According to a classical result of Jamison [9], an almost cover of the $n$-dimensional affine space over the $q$-element finite field requires at least $(q-1)n$ hyperplanes. Equivalently, to pierce every affine hyperplane in $\mathbb{F}_{q}^{n}$ one needs at least $(q-1)n+1$ points, see [6]. See also [5] for further results in finite geometries. Another example is the Alon–Füredi theorem [3]: Every almost cover of the vertex set of an $n$-dimensional cube requires at least $n$ hyperplanes. The key observation of the present note is the following analogue.
Theorem 1.
Every almost cover of the vertices of $\Pi_{n-1}$ consists of at least ($n\choose 2$ hyperplanes. This bound is sharp.
In comparison, all vertices of $\Pi_{n-1}$ are contained in just one hyperplane, and even when $\Pi_{n-1}$ is embedded into $\mathbb{R}^{n-1}$, its vertices can be covered by $n$ hyperplanes, see Section 3. Our proof of the above theorem relying on the polynomial method parallels that of the Alon–Füredi theorem. After recalling the Combinatorial Nullstellensatz, it is given in Section 3. In Section 4 we extract a version of the polynomial lemma that may be useful in the study of permutations. We conclude this paper by suggesting a common generalization of Theorem 1 and the Alon–Füredi theorem to zonotopes.
2. Preliminaries
The Combinatorial Nullstellensatz, formulated by Noga Alon in the late nineties, describes, in an efficient way, the structure of multivariate polynomials whose zero-set includes a Cartesian product over a field $\mathbb{F}$. This characterization immediately implies the following non-vanishing criterion, see [1].
Theorem 2.
Let $f=f(x_{1},\ldots,x_{n})$ be a polynomial in $\mathbb{F}[x_{1},\ldots,x_{n}]$. Suppose that there is a monomial $\prod_{i=1}^{n}x_{i}^{d_{i}}$ such that $\sum_{i=1}^{n}d_{i}$ equals the degree of $f$ and whose coefficient in $f$ is nonzero. If $S_{1},\ldots,S_{n}$ are subsets of $\mathbb{F}$ with $|S_{i}|>d_{i}$, then there are $s_{1}\in S_{1},\ldots,s_{n}\in S_{n}$ such that $f(s_{1},\ldots,s_{n})\neq 0$.
Based on this one can prove Theorem 1 via Theorem 7 derived in Section 4. To present a more transparent direct proof we recall an immediate consequence of Theorem 5 in [3].
Theorem 3.
Let $S_{1},\dots,S_{n}$ be nonempty subsets of $\mathbb{F}$, $B=S_{1}\times\dots\times S_{n}$. If a polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ vanishes at every point of $B$ except one, then its degree is at least $\sum_{i=1}^{n}(|S_{i}|-1)$.
Remark. This result can be derived from Theorem 2 as well. It can also be viewed as a special case of Corollary 4.2 in [4]. Whichever tool one applies the background is the same: reduction modulo a natural Gröbner basis coupled with Lagrange interpolation.
3. Covers of the permutohedron
All points of $P_{n}$ lie in the hyperplane $H_{0}$ of equation $f_{0}(x)=\sum_{i=1}^{n}x_{i}-{{n+1}\choose 2}=0$. However, if we view $\Pi_{n-1}$ as an $(n-1)$-dimensional polytope, the meaningful question is the following: How many affine hyperplanes different from $H_{0}$ are needed to cover all points of $P_{n}$? Surprisingly this appears to be a more elusive problem. It is clear that the hyperplanes of equation $x_{1}=i$, $1\leq i\leq n$ cover $P_{n}$. Each of them contains exactly $(n-1)!$ points of $P_{n}$, and their intersections with $P_{n}$ are pairwise disjoint. The same is true for the hyperplanes of equation $x_{i}=1$. If $n\geq 4$ is even, then one can do even better: The hyperplanes of equation $x_{1}+x_{j}=n+1$, $2\leq j\leq n$ cover $P_{n}$. Each of them contains exactly $n(n-2)!$ points of $P_{n}$, and their intersections with $P_{n}$ are pairwise disjoint. We believe that these examples are extremal in the following sense.
Conjecture 4.
Suppose that the vertex set of $\Pi_{n-1}$ is contained in the union of the hyperplanes $H_{1},\ldots,H_{m}$ different from $H_{0}$. If $n$ is odd, then $m\geq n$. If $n\geq 4$ is even, then $m\geq n-1$.
Consider a hyperplane $H$ not parallel to $H_{0}$, it intersects $H_{0}$ in a 1-codimensional affine subspace. To cover $P_{n}$ by $n$ or less such hyperplanes one needs to find such an $H$ which intersects $P_{n}$ in at least $(n-1)!$ points. $H$ has an equation of the form $f(x)=0$, where $f$ is a linear polynomial, in this case we write $H=H(f)$. Then $H(g)\cap H_{0}=H(f)\cap H_{0}$ if and only if there exist $\alpha,\beta\in\mathbb{R}$, $\alpha\neq 0$ such that $g=\alpha f+\beta f_{0}$. Apart from such equivalence, it seems that $H(f)$ intersects $P_{n}$ in more than $(n-1)!$ points if and only if $n$ is even and $f=x_{i}+x_{j}-(n+1)$ for some $i\neq j$. $|H(f)\cap P_{n}|=(n-1)!$ occurs in each dimension for $f=x_{i}-x_{j}-1$, and also for $f=x_{i}+x_{j}-n$, $f=x_{i}+x_{j}-(n+1)$, $f=x_{i}+x_{j}-(n+2)$ when $n$ is odd. From these examples one can construct various economical hyperplane covers of $P_{n}$. For example, the hyperplanes of equation $x_{n}=1$, $x_{n}-x_{i}=1$, $11\leq i\leq n-$ 1 cover $P_{n}$ for every $n$, whereas the hyperplanes of equation $x_{1}=(n+1)/2$, $x_{1}+x_{j}=n+1$, $2\leq j\leq n$ cover $P_{n}$ when $n$ is odd. We believe that the following is true.
Conjecture 5.
If $n$ is odd, then every hyperplane different from $H_{0}$ contains at most $(n-1)!$ points of $P_{n}$.
If true, this would imply Conjecture 4 as follows. The statement follows immediately if $n$ is odd. Suppose that $n\geq 4$ is even and the hyperplanes $H_{1},\dots,H_{m}$ cover $P_{n}$. We may assume that $m\leq n-1$. Consider the hyperplanes $H(x_{j}-n)$ for $1\leq j\leq n$. They intersect $\Pi_{n-1}$ in pairwise disjoint respective facets $F_{j}$, congruent to $\Pi_{n-2}$ apiece, of which no two can be contained in the same hyperplane $H_{i}$. By the pigeonhole principle there is a facet $F_{j}$, which is not fully contained in any $H_{i}$, $1\leq i\leq m$. Within the $(n-1)$-dimensional ambient space $H_{0}$, consider the $k\leq m$ hyperplanes $H_{i}\cap H_{0}$ which intersect the affine hull of $F_{j}$. They cover all the vertices of $F_{j}$ and are different from its affine hull $H(x_{j}-1)\cap H_{0}$. Applying a suitable isometry, it follows from Conjecture 5 that each of them contains at most $(n-2)!$ vertices of $F_{j}$. Accordingly, $k(n-2)!\geq(n-1)!$, yielding $m\geq k\geq n-1$.
Turning to almost covers of the vertex set of $\Pi_{n-1}$, note that unlike to the case of exact covers it does not matter, which ambient space we consider.
Proof of Theorem 1. A key observation, due to Alon [2], which leads to a by now standard application of the Combinatorial Nullstellensatz is that the permutations of an $n$-element set $A$ can be identified with the set obtained from $A^{n}$ by removing the roots of the Vandermonde polynomial $V({x})=\prod_{i<j}(x_{j}-x_{i})$. Geometrically, $P_{n}$ is obtained from $\{1,2,\ldots,n\}^{n}$ by removing the points covered by the ($n\choose 2$ hyperpanes of equation $x_{i}=x_{j}$.
Suppose that a collection of hyperplanes defined by the linear equations $f_{1}({x})=0,\dots,f_{m}({x})=0$ constitute an almost cover of $P_{n}$ where the uncovered vertex is ${v}$. Then the polynomial
$f(x)=V({x})\prod_{i=1}^{m}f_{i}({x})\in\mathbb{R}[x_{1},\ldots,x_{n}]$
of degree (${n\choose 2}+m$ attains the value 0 at every point of the Cartesian product $\{1,2,\dots,n\}^{n}$ except ${v}$. According to Theorem 3,
(${n\choose 2}+m\geq n(|A|-1)=n(n-1),$
hence $m\geq{n\choose 2}$. To see that this bound cannot be improved, notice that the hyperplanes $x_{i}=j$ ($i<j$) cover every vertex but $(1,2,\ldots,n)$. ∎
Remarks. Replacing $\prod f_{i}$ by an arbitrary polynomial of degree $m$ in the above proof one obtains the following more general theorem: If an algebraic surface covers all vertices of $\Pi_{n-1}$ except one, then its degree is at least ($n\choose 2$. The proof also works verbatim for sets of the form
$X(\alpha_{1},\ldots,\alpha_{n})=\{(\alpha_{\pi(1)},\ldots,\alpha_{\pi(n)})\mid\pi\in S_{n}\},$
where $\alpha_{1},\ldots,\alpha_{n}$ are different elements of an arbitrary field $\mathbb{F}$.
Corollary 6.
Let $v,u_{1},\dots,u_{k}\in X_{n}=X(\alpha_{1},\ldots,\alpha_{n})$ such that $v$ is not contained in the affine hull of $\{u_{1},\dots,u_{k}\}$. Suppose that the affine hyperplanes $H_{1},\dots,H_{m}$ satisfy
$X_{n}\setminus\{u_{1},\dots,u_{k}\}\subseteq H_{1}\cup\dots\cup H_{m},\quad v\not\in H_{1}\cup\dots\cup H_{m}.$
Then $m\geq{n\choose 2}-1$.
Proof.
Choose an affine hyperplane $H$ which contains the affine hull of $\{u_{1},\dots,u_{k}\}$ but does not contain $v$. Then $H,H_{1},\dots,H_{m}$ constitute an almost cover of $X_{n}$, hence $m+1\geq{n\choose 2}$. ∎
4. Permutations via polynomials
Besides Theorem 1 and Snevily’s problem treated in [2, 7], there are several combinatorial problems regarding permutations that may be treated with the polynomial method, see e.g. [8, 10]. Although it is not needed for the present work, it still may be useful to formulate the following variant of Theorem 2 for permutations, which can be extracted from the proofs in [2, 7].
Theorem 7.
Let $\alpha_{1},\ldots,\alpha_{n}$ be $n$ different elements of a field $\mathbb{F}$. Let $f\in\mathbb{F}[x]$ be a polynomial of degree (${n\choose 2}$. For each permutation $\pi\in S_{n}$, denote by $c_{\pi}$ the coefficient of the monomial $\prod_{i=1}^{n}x_{i}^{\pi(i)-1}$ in $f$. If
$\sum_{\pi\in S_{n}}\mbox{\rm sgn}(\pi)c_{\pi}\neq 0,$
then there exists a point $s\in X(\alpha_{1},\ldots,\alpha_{n})$ such that $f(s)\neq 0$.
Proof.
Put $S_{i}=\{\alpha_{1},\ldots,\alpha_{n}\}$ and consider the polynomial $g(x)=V(x)f(x)$, then $\deg g=n(n-1)$. Using the determinant expression
$V(x)=\mbox{\rm Det}(x_{i}^{j-1})_{1\leq i,j\leq n}=\sum_{\pi\in S_{n}}\mbox{\rm sgn}(\pi)\prod_{i=1}^{n}x_{i}^{\pi(i)-1}$
it is easy to compute the coefficient $c$ of the monomial $\prod_{i=1}^{n}x_{i}^{n-1}$ in $g$. For each $\pi\in S_{n}$ denote by $\pi^{\prime}$ the permutation satisfying $\pi^{\prime}(i)=n+1-\pi(i)$ for every $i$, then $(\pi^{\prime})^{\prime}=\pi$ and $\mbox{\rm sgn}(\pi^{\prime})=(-1)^{n\choose 2}\mbox{\rm sgn}(\pi)$. Thus,
$c=\sum_{\pi\in S_{n}}\mbox{\rm sgn}(\pi)c_{\pi^{\prime}}=\sum_{\pi^{\prime}\in S_{n}}\mbox{\rm sgn}(\pi^{\prime})c_{(\pi^{\prime})^{\prime}}=(-1)^{n\choose 2}\sum_{\pi\in S_{n}}\mbox{\rm sgn}(\pi)c_{\pi}\neq 0.$
Since $g(x)=0$ for $x\in(S_{1}\times\dots\times S_{n})\setminus X(\alpha_{1},\ldots,\alpha_{n})$, the result follows from Thm 2. ∎
As a quick application we derive Lemma 4 in [7]. For an $n\times n$ matrix $M=(m_{ij})$, the permanent of $M$ is $\mbox{\rm Per}M=\sum_{\pi\in S_{n}}m_{1\pi(1)}m_{2\pi(2)}\ldots m_{n\pi(n)}$.
Corollary 8.
Let $a_{1},a_{2},\ldots,a_{n}\in\mathbb{F}$ such that $\mbox{\rm Per}V(a_{1},\ldots,a_{n})\neq 0$. Then, for any subset $B\subset\mathbb{F}$ of cardinality $n$ there is a numbering $b_{1},\ldots,b_{n}$ of the elements of $B$ such that the products $a_{1}b_{1},\ldots,a_{n}b_{n}$ are pairwise different.
Proof.
Take $f(x)=V(a_{1}x_{1},\dots,a_{n}x_{n})=\prod_{1\leq i<j\leq n}(a_{j}x_{j}-a_{i}x_{i})$, then
$f(x)=\sum_{\pi\in S_{n}}\mbox{\rm sgn}(\pi)\prod_{i=1}^{n}(a_{i}x_{i})^{\pi(i)-1}.$
Thus, $c_{\pi}=\mbox{\rm sgn}(\pi)\prod_{i=1}^{n}a_{i}^{\pi(i)-1}$ and
$\sum_{\pi\in S_{n}}\mbox{\rm sgn}(\pi)c_{\pi}=\sum_{\pi\in S_{n}}(\mbox{\rm sgn}(\pi))^{2}\prod_{i=1}^{n}a_{i}^{\pi(i)-1}=\mbox{\rm Per}V(a_{1},\ldots,a_{n})\neq 0.$
The claim follows applying Theorem 7 with $\{\alpha_{1},\ldots,\alpha_{n}\}=B$. ∎
5. Concluding remarks
For a finite set of points $X$ in an affine space and $v\in X$ define $\mbox{\rm ac}(X,v)$ as the minimum number of hyperplanes whose union contains $X\setminus\{v\}$ but does not contain $v$ and let
$\mbox{\rm ac}(X)=\min_{v\in X}\mbox{\rm ac}(X,v).$
If $X$ is the vertex set of a convex polytope, we simply write $\mbox{\rm ac}(P)$. Thus, $\mbox{\rm ac}(C_{n})=n$, and $\mbox{\rm ac}(\Pi_{n-1})={n\choose 2}$. For $n$-element sets $X\subset\mathbb{R}^{d}$, $\mbox{\rm ac}(X)$ attains every integer between $1$ and $\lceil(n-1)/d\rceil$, even when $X$ is restricted to sets in convex position. Thus one may expect a meaningful estimate on $\mbox{\rm ac}(X)$ only if $X$ satisfies some appropriate geometric, combinatorial or algebraic condition. Inspired by Theorem 1 we found that being the vertex set of a zonotope may be a reasonable indicator.
A zonotope is a convex polytope that can be represented as the Minkowski sum of a finite number of line segments. A collection of line segments is called nondegenerate if no two of the segments are parallel to each other. Each zonotope $Z$ can be written as the Minkowski sum of a nondegenerate collection of line segments. Such a representation is unique up to translations of the segments by vectors which add up to 00{0}. The number of the summands, denoted by $\mbox{\rm rk}(Z)$, we call the rank of $Z$.
Each zonotope of rank $n$ can be obtained as an orthogonal projection of an affine image of the $n$-dimensional unit cube $C_{n}$, see [12]. If the dimension of a zonotope is 1, then it is a line segment. The 2-dimensional zonotopes of rank $n$ are exactly the centrally symmetric convex $2n$-gons, and every almost cover of such a polygon with lines requires at least $n$ lines. These are the cases when most vertices of $C_{n}$ are lost along the projection. In the other extreme, when the projection is the identical transformation, $\mbox{\rm ac}(Z)=\mbox{\rm rk}(Z)$ follows from the Alon–Füredi theorem, and Theorem 1 asserts the same for $Z=\Pi_{n-1}$.
Conjecture 9.
Every almost cover of the vertices of a zonotope $Z$ consists of at least $\mbox{\rm rk}(Z)$ hyperplanes.
Further examples supporting this conjecture can be found in [11], where we also investigate the extent to which our method can be applied to Coxeter permutahedra. If true, this bound would be best possible in the following sense. Suppose that the full dimensional zonotope $Z$ is the Minkowski sum of the line segments $[0,v_{i}]$, $1\leq i\leq n$. Denoting by $H$ the affine hull of the points $v_{i}$ and assuming $0\not\in H$, the hyperplanes $H,2H,\dots,nH$ cover each vertex of $Z$ except 00. Since an arbitrary scaling of the generating vectors of a zonotope preserves the face lattice of the zonotope, each zonotope is combinatorially equivalent to a zonotope $Z$ which has an almost cover with $\mbox{\rm rk}(Z)$ hyperplanes.